OpenStudy (anonymous):

given that 6(y+3) = 2(x+9) , find the ratio of x to y.

OpenStudy (anonymous):

i think is 3/1

OpenStudy (anonymous):

First expand the eqn, collect like terms and then write x and y in fraction form

OpenStudy (anonymous):

6 (y+3) = 2 (x+9) 3 (y+3) = (x+9) 3y + 9 = x + 9 3y = x Ratio: y to x = 3 to 1, or x to y is 1 to 3.

OpenStudy (anonymous):

Good job! :)

OpenStudy (anonymous):

Ratio is 1:3

OpenStudy (anonymous):

And not 3:1